# Groups Acting Acylindrically on Trees

### v5

@@ 3.2. Acylindrical Hyperbolicity of Graph Products
Let $G$ be the RAAG on the graph $\Gamma$ shown to the left of Figure 2. First we will show that $G$ is acylindrically hyperbolic, and by Corollary 3.11 it suffices to show that $\Gamma$ is irreducible and that $G$ is not virtually cyclic. The second of these conditions is clear as $G$ has a $\mathbb{Z}^{2}$ subgroup, and for the first, the graph theoretical complement of $\Gamma$ is the connected graph shown in Figure 2 (right), so $\Gamma$ is irreducible and $G$ is acylindrically hyperbolic as claimed.
Next we will show that $G$ does not have property (FA-), and in fact has no infinite normal subgroups with (FA-). A result of Antolin and Minasyan asserts that any subgroup of a right angled Artin group is either free abelian of finite rank or projects onto a free group of rank two [Antolinminasyan11, Corollary 1.6]. Thus $G$ does not have (FA-) as it is not free abelian, and any non-trivial (and thus infinite) normal subgroup of $G$ with property (FA-) would have to be free abelian as a projection to the free group of rank two would induce an action on a tree with no fixed points or line. However, the fact that $G$ is acylindrically hyperbolic implies that $G$ can contain no infinite normal (or indeed even s-normal) abelian subgroup [Osin13, Lemma 7.1]. Thus $G$ has no normal subgroup with property (FA-) as claimed.
Finally $G$ is not acylindrically arboreal by Corollary 3.12 as $\operatorname{diam}(\Gamma)=2$, so $G$ has all properties required. ∎
@@
We can also consider subgroups of graph products using Theorem 1.3. Recall that for a graph product of groups $G=\mathbb{GP}(\Gamma,\mathcal{G})$ and a subgroup $H$ of $G$, the parabolic closure of $\operatorname{Pc_{\Gamma}}_{\Gamma}(H)$ to be the intersection of all parabolic subgroups of $G$ that contain $H$, and the essential support $\operatorname{esupp}_{\Gamma}(H)$ to be the essential support of the parabolic closure of $H$.
@@
Let $G$ be the the RAAG associated on a $P_{3}$ whose vertices groups are generated by the letters $a,b$ and $c$ so that $a$ and $c$ generate a copy of $F_{2}$. The subgroup $\langle ab,bc\rangle$ is isomorphic to $F_{2}$ so is acylindrically arboreal, but its essential support is $\Gamma$, which contains no separated pair of vertices.
@@
In this section we will consider the acylindrical arboreality of certain fundamental groups of 3-manifold, and provide a proof of Theorem 1.8. We will provide a geometric condition on the fundamental group of a compact and orientable hyperbolic 3-manifold with empty or incompressible toroidal boundary which is equivalent to acylindrical arboreality, where by hyperbolic we mean carries a complete metric of constant negative curvature on its interior. For avoidance of doubt, all surface embeddings will be assumed to be proper, i.e. if $f:S\hookrightarrow M$ is an embedding of a surface into a $3$-manifold $M$ then $f^{-1}(\partial M)=\partial S$.
@@ 4.1. Relative Quasi-Convexity and Malnormality
The concept of malnormality is very closely linked with acylindricity, and in the context of relatively hyperbolic groups this is very well studied, but it can be difficult to find malnormal subgroups. We will instead use the related concept of relative quasi-convexity, for which we will use the definition due to Osin [Osin06, Definition 1.8]. For a more thorough treatment of relative hyperbolicity and relative quasi-convexity, see [Hruska08].
@@
Let $G$ be a group hyperbolic relative to a collection of subgroups $\mathcal{P}$ such that $G$ is generated by the finite set $X$. We say that a subgroup $H$ of $G$ is relatively quasi-convex with respect to $\mathcal{P}$, or quasi-convex relative to $\mathcal{P}$, if there exists some constant $\epsilon>0$ such that the following condition holds. Let $f$ and $g$ be elements of $H$, and $p$ a geodesic between $f$ and $g$ in the Cayley graph of $G$ with respect to the generating set $X\cup\mathcal{P}$. Then for any vertex $w\in p$ there exists a vertex $v\in H$ such that
$d_{X\cup\mathcal{P}}(v,w)\leq\epsilon.$
We have the following theorem that links relative quasi-convexity to malnormality.
@@
[HW09, Theorem 1.4] Let $G$ be a group hyperbolic relative to a collection of subgroups $\mathcal{P}$, and $H$ a relatively quasi-convex subgroup of $G$ with respect to $\mathcal{P}$. Then there exists some constant $n_{H}\in\mathbb{Z}_{>0}$ such that for any set $g_{1}H,...,g_{n_{H}}H$ of $n_{H}$ distinct cosets of $H$ in $G$ the intersection ${\bigcap_{0<i\leq n_{H}}g_{i}Hg_{i}^{-1}}$ is either finite or is a peripheral subgroup of $G$.
We call a set of conjugates that arise from disjoint cosets in this way essentially disjoint. A special case of relative quasi-convexity occurs when $H$ is a subgroup of a hyperbolic group $G$ (i.e. $\mathcal{P}$ contains only the trivial subgroup). In this case we say that the subgroup in question is quasi-convex. The restriction of Theorem 4.2 to quasi-convex subgroups of hyperbolic groups was first proved in [GitikMahanRipsSageev98, Main Theorem], and gives the stronger conclusion that the intersection of a certain number of cosets of $H$ must be finite, as we have no non-trivial peripheral elements. This implies that the action of a group on a tree with quasi-convex edge stabilisers and finitely many orbits of edges will be acylindrical, a fact that is well known among experts (see [Linton22], for example).
@@ 4.2. Splittings of 3-Manifold Groups and Subgroup Tameness
We will require some deep results from various authors in the field of 3-manifold groups. We begin with an important definition.
Let $M$ be a compact and orientable 3-manifold and $S$ a compact surface embedded in $M$ such that $S\not\cong S^{2}$. A disk $D\subset M$ is a compressing disk for $S$ if $S\cap D=\partial D$ and this intersection is transverse. A compressing disk is non-trivial if $\partial D$ does not bound a disk in $S$. We say that $S$ is compressible if it admits a non-trivial compressing disk, and we say that $S$ is incompressible otherwise. By convention, a sphere $S^{2}$ embedded in a $3$-manifold $M$ is considered to be incompressible if it does not bound a ball, and a disc embedded in a $3$-manifold is incompressible if it is not homotopic into a boundary component. We say that a $3$-manifold $M$ is irreducible if every sphere embedded in $M$ bounds a $3$-ball, or equivalently if $M$ contains no embedded incompressible sphere.
An incompressible surface $S\subset M$ is called $2$-sided if there exists an embedding $h\colon S\times[-1,1]\rightarrow M$ such that $h(x,0)=x$ for all $x\in S$. We have the following lemma.
@@
[Hempel76, Corollary 6.2] Let $M$ be a $3$-manifold and $S$ a $2$-sided incompressible surface in $M$. Then the natural map $i_{*}\colon\pi_{1}(S)\rightarrow\pi_{1}(M)$ induced by the inclusion $i\colon S\rightarrow M$ is injective.
@@
We say that a $3$-manifold $M$ is fibred if it admits the structure of a surface bundle over $S^{1}$. In such a case there exists a $2$-sided incompressible surface $S\subset M$ such that $\pi_{1}(M)\cong\pi_{1}(S)\rtimes\mathbb{Z}$. We say that a subgroup $H\leq\pi_{1}(M)$ is a surface fibre subgroup of $\pi_{1}(M)$ if it is finitely generated and if $M$ is admits a surface bundle structure over $S^{1}$ with fibre surface $S$ such that $H$ is the image of $\pi_{1}(S)$ in $\pi_{1}(M)$.
We say that $H$ is a virtual surface fibre subgroup of $\pi_{1}(M)$ if there exists some finite index subgroup $H^{\prime}\leq\pi_{1}(M)$ such that $H\leq H^{\prime}$ and such that $H$ is a fibre subgroup of $H^{\prime}$. In this case $M$ has a finite sheeted cover whose fundamental group contains $H$ and in which $H$ is the fundamental group of a fibre.
@@
[Stallings1961OnFC, Theorems 1, 2] Let M be a compact and irreducible 3-manifold and let $H\unlhd\pi_{1}(M)$ be a finitely generated normal subgroup. If $\pi_{1}(M)/H\cong\mathbb{Z}$ and $H\not\cong\mathbb{Z}/2\mathbb{Z}$, then $M$ can be expressed as a fibred 3-manifold with fibre surface $S$. Furthermore, we can choose $S$ such that $\pi_{1}(S)=H$
@@
[Scott72, Theorem 2.3] Let $M$ be a closed and orientable irreducible 3-manifold, and suppose $\pi_{1}(M)\cong A*_{C}B$ where $C\neq A$ or $B$ and $C$ is isomorphic to the fundamental group of a closed surface $S$. Then there is an incompressible embedding of $S$ in $M$ separating $M$ into $M_{1}$ and $M_{2}$ with $\pi_{1}(M_{1})=A$, $\pi_{1}(M_{2})=B$ and $\pi_{1}(S)=C$.
More generally, have the following powerful result essentially due to Stallings, Epstein and Waldhausen that can be used to construct 2-sided incompressible surfaces from any graph of groups decomposition of the fundamental group of a closed and orientable 3-manifold. We will use the version that appears in [Culler1983VarietiesOG].
@@
[Culler1983VarietiesOG, Proposition 2.3.1] Let $M$ be a compact orientable 3-manifold. For any non-trivial graph of groups decomposition $(\Gamma,\mathfrak{G})$ of $\pi_{1}(M)$ with finitely many edge groups there exists a non-empty system $\overline{\Sigma}=\{\Sigma_{1},...,\Sigma_{n}\}$ of compact $2$-sided incompressible surfaces embedded in $M$, none of which are boundary parallel, such that for all $i$, $\operatorname{Im}(\pi_{1}(\Sigma_{i})\rightarrow\pi_{1}(M))$ is contained in some edge group of $(\Gamma,\mathfrak{G})$ and for all connected components $M_{j}$ of $M\backslash\overline{\Sigma}$, $\operatorname{Im}(\pi_{1}(M_{j})\rightarrow\pi_{1}(M))$ is contained in some vertex group of $(\Gamma,\mathfrak{G})$.
@@
In the paper of Culler and Shalen, the authors do not conclude that the surfaces they construct are compact. However, compactness is an immediate consequence of their proof — the surfaces are constructed as the connected components of the continuos pre-image in the compact manifold of a closed set, so are themselves compact.
@@
[AschenbrennerFriedlWilton12, (K.18)] Let $M$ be a compact and orientable hyperbolic $3$-manifold with empty or toroidal boundary. Then $\pi_{1}(M)$ is hyperbolic relative to the conjugacy classes of the subgroups arising from the boundary components of $M$.
@@
Let $M$ be a compact and orientable hyperbolic manifold with empty or toroidal boundary. Then $M$ is irreducible.
@@
Let $M$ be as in the statement. Then the universal cover of the interior of $M$ is a copy of $\mathbb{H}^{3}$, which has vanishing $\pi_{2}$. It follows that $\pi_{2}$ of the interior of $M$ must also vanish, and the boundary of $M$ contains no spherical components, so $M$ is irreducible as required as it can contain no non-trivial sphere. ∎
@@
A compact and orientable hyperbolic manifold $M$ with empty or toroidal boundary that contains an embedded $2$-sided incompressible disc must be a solid torus. In particular, as a consequence of Theorem 4.6, $\pi_{1}(M)$ admits a non-trivial splitting over a trivial subgroup if and only if $M$ is a solid torus.
@@
For the first part of this corollary, assume that $M$ contains an embedded disc , so there exists an embedding of a disc $D$ into $M$ such that the boundary of $D$ is a non-trivial curve embedded in some boundary component $T$ of $M$. We will argue as in [HatcherNotes, Page 14, (3)] that $M$ is a solid torus.
@@
[AschenbrennerFriedlWilton12, (C.2)] Let $M$ be an orientable, irreducible $3$-manifold with empty or toroidal boundary and infinite fundamental group. Then $\pi_{1}(M)$ is torsion free.
Many of our methods will rely on a powerful result known as subgroup tameness to classify subgroups of the fundamental group of hyperbolic manifolds. For full a definition of the concept of geometric finiteness we direct the reader to Chapter 5 of the book 3-Manifold Groups by Aschenbrenner, Friedl and Wilton [AschenbrennerFriedlWilton12]. For our purposes however, we will be able to define geometrically finite subgroups simply as relatively quasi-convex subgroups of hyperbolic $3$-manifold groups using the following theorem of Hruska.
@@
[Hruska08, Corollary 1.6] Let $G=\pi_{1}(M)$ be the fundamental group of a compact and orientable hyperbolic 3-manifold with empty or toroidal boundary and let $H$ be a finitely generated subgroup of G. Then $H$ is geometrically finite in $G$ if and only if $H$ is relatively quasi-convex with respect to the subgroups of $G$ the correspond to the boundary components.
@@
[AschenbrennerFriedlWilton12, Theorem 5.2] Let $M$ be a hyperbolic 3-manifold and let $H\leq\pi_{1}(M)$ be a finitely generated subgroup. Then either

### v6

@@ 3.2. Acylindrical Hyperbolicity of Graph Products
Let $G$ be the RAAG on the graph $\Gamma$ shown to the left of Figure 2. First we observe that $G$ is not acylindrically arboreal by Corollary 3.12 as $\operatorname{diam}(\Gamma)=2$.
Next we show that $G$ is acylindrically hyperbolic, and by Corollary 3.11 it suffices to show that $\Gamma$ is irreducible and that $G$ is not virtually cyclic. The second of these conditions is clear as $G$ has a $\mathbb{Z}^{2}$ subgroup, and for the first, the graph theoretical complement of $\Gamma$ is the connected graph shown in Figure 2 (right), so $\Gamma$ is irreducible and $G$ is acylindrically hyperbolic as claimed.
Finally we show that $G$ does not have property (FA-), and in fact has no infinite normal subgroups with (FA-). A result of Antolín and Minasyan asserts that any subgroup of a right angled Artin group is either free abelian of finite rank or projects onto a free group of rank two [undef, Corollary 1.6]. Thus $G$ does not have (FA-) as it is not free abelian. Further, any non-trivial (and thus infinite) normal subgroup of $G$ with property (FA-) would have to be free abelian as a projection to the free group of rank two would induce an action on a tree with no fixed points or line. However, the fact that $G$ is acylindrically hyperbolic implies that $G$ can contain no infinite normal (or indeed even s-normal) abelian subgroup [undefaf, Lemma 7.1]. Thus $G$ has no normal subgroup with property (FA-) as claimed, and so $G$ has all properties required. ∎
@@
We can also consider subgroups of graph products using Theorem 1.3. Recall that, for a graph product of groups $G=\mathbb{GP}(\Gamma,\mathcal{G})$ and a subgroup $H$ of $G$, the parabolic closure $\operatorname{Pc_{\Gamma}}(H)$ is the intersection of all parabolic subgroups of $G$ that contain $H$, and the essential support $\operatorname{esupp}_{\Gamma}(H)$ is the essential support of the parabolic closure of $H$.
@@
Let $G$ be the RAAG on a $P_{3}$ whose vertices groups are generated by the letters $a,b$ and $c$ so that $a$ and $c$ generate a copy of $F_{2}$. The subgroup $\langle ab,bc\rangle$ is isomorphic to $F_{2}$ so is acylindrically arboreal, but its essential support is $\Gamma$, which contains no separated pair of vertices.
@@
In this section we will consider the acylindrical arboreality of certain fundamental groups of 3-manifolds, and provide a proof of Theorem 1.8. We will provide a geometric condition on the fundamental group of a compact and orientable hyperbolic 3-manifold with empty or incompressible toroidal boundary which is equivalent to acylindrical arboreality, where by hyperbolic we mean carries a complete metric of constant negative curvature on its interior. For avoidance of doubt, all surface embeddings will be assumed to be proper, i.e. if $f:S\hookrightarrow M$ is an embedding of a surface into a $3$-manifold $M$ then $f^{-1}(\partial M)=\partial S$.
@@ 4.1. Relative Quasi-Convexity and Malnormality
Let $G$ be a group, and $H$ a subgroup of $G$. We say that $H$ is malnormal in $G$ if for all $g\in G\setminus H$, $gHg^{-1}\cap H=\{1\}$. The concept of malnormality is very closely linked with acylindricity, and in the context of relatively hyperbolic groups this is very well studied, but it can be difficult to find malnormal subgroups. We will instead use the related concept of relative quasi-convexity, for which we will use the definition due to Osin [undefah, Definition 1.8]. For a more thorough treatment of relative hyperbolicity and relative quasi-convexity, see [undefs].
@@
Let $G$ be a group hyperbolic relative to a collection of subgroups $\mathcal{P}$ such that $G$ is generated by the finite set $X$. We say that a subgroup $H$ of $G$ is relatively quasi-convex with respect to $\mathcal{P}$, or quasi-convex relative to $\mathcal{P}$, if there exists some constant $\epsilon>0$ such that the following condition holds. Let $f$ and $g$ be elements of $H$, and $p$ a geodesic between $f$ and $g$ in the Cayley graph of $G$ with respect to the generating set $X\cup\mathcal{P}$. Then for any vertex $w\in p$ there exists a vertex $v\in H$ such that $d_{X\cup\mathcal{P}}(v,w)\leq\epsilon$.
We have the following definitions and theorem that link relative quasi-convexity to malnormality.
Definition 4.2.
Let $G$ be a group and let $H$ be a subgroup of $G$. Let $g_{1},...,g_{n}\in G$ be a set of $n$ elements in $G$. If $g_{1}H,...,g_{n}H$ is a set of disjoint cosets of $H$ in $G$, we say that the conjugates $g_{1}Hg_{1}^{-1},...,g_{n}Hg_{n}^{-1}$ of $H$ in $G$ are essentially distinct.
Now assume that $G$ is hyperbolic relative to a collection of subgroups $\mathcal{P}$. We say that a subgroup $H$ of $G$ has finite relative height (in $G$) if there exists some constant $n_{H}\in\mathbb{Z}_{>0}$ such that for any set $g_{1}Hg_{1}^{-1},...,g_{n_{H}}Hg_{n_{H}}^{-1}$ of $n_{H}$ essentially distinct conjugates of $H$ in $G$ we have that ${\bigcap_{0<i\leq n_{H}}g_{i}Hg_{i}^{-1}}$ is either finite or is a peripheral subgroup of $G$.
@@
[undeft, Theorem 1.4] Let $G$ be a group hyperbolic relative to a collection of subgroups $\mathcal{P}$, and $H$ a relatively quasi-convex subgroup of $G$ with respect to $\mathcal{P}$. Then $H$ has finite relative height in $G$.
A special case of relative quasi-convexity occurs when $H$ is a subgroup of a hyperbolic group $G$ (i.e. $\mathcal{P}$ contains only the trivial subgroup). In this case we say that the subgroup in question is quasi-convex. The restriction of Theorem 4.3 to quasi-convex subgroups of hyperbolic groups was first proved in [undefl, Main Theorem], and gives the stronger conclusion that the intersection of a certain number of cosets of $H$ must be finite, as we have no non-trivial peripheral elements. This implies that the action of a hyperbolic group on a tree with quasi-convex edge stabilisers and finitely many orbits of edges will be acylindrical, a fact that is well known among experts (see [undefy], for example).
@@ 4.2. Splittings of 3-Manifold Groups and Subgroup Tameness
We will require some deep results from various authors in the field of 3-manifold groups. We begin with some important definitions.
Definition 4.4.
Let $M$ be a compact and orientable 3-manifold and $S$ a compact surface embedded in $M$.
  1. (1)
    We say that $S$ is boundary parallel if there exists an isotopy of $S$ onto some boundary component of $M$.
  2. (2)
    Assume $S\not\cong S^{2}$. A disc $D\subset M$ is a compressing disc for $S$ if $S\cap D=\partial D$ and this intersection is transverse. A compressing disc is non-trivial if $\partial D$ does not bound a disc in $S$. We say that $S$ is compressible if it admits a non-trivial compressing disc, and we say that $S$ is incompressible otherwise.
    By convention, a sphere $S^{2}$ embedded in a $3$-manifold $M$ is considered to be incompressible if it does not bound a ball, and a disc embedded in a $3$-manifold is incompressible if it is not homotopic into a boundary component.
  3. (3)
    We say that $M$ is irreducible if every sphere embedded in $M$ bounds a $3$-ball, or equivalently if $M$ contains no embedded incompressible sphere.
  4. (4)
    If $S$ is incompressible, we say that $S$ is $2$-sided if there exists an embedding $h\colon S\times[-1,1]\rightarrow M$ such that $h(x,0)=x$ for all $x\in S$.
We then have the following lemma.
@@
[undefr, Corollary 6.2] Let $M$ be a $3$-manifold and $S$ a $2$-sided incompressible surface in $M$. Then the natural map $i_{*}\colon\pi_{1}(S)\rightarrow\pi_{1}(M)$ induced by the inclusion $i\colon S\rightarrow M$ is injective.
@@
We say that a $3$-manifold $M$ is fibred if it admits the structure of a surface bundle over $S^{1}$. In such a case there exists a $2$-sided incompressible surface $S\subset M$ such that $\pi_{1}(M)\cong\pi_{1}(S)\rtimes\mathbb{Z}$. We say that a subgroup $H\leq\pi_{1}(M)$ is a surface fibre subgroup of $\pi_{1}(M)$ if it is finitely generated and if $M$ admits a surface bundle structure over $S^{1}$ with fibre surface $S$ such that $H$ is the image of $\pi_{1}(S)$ in $\pi_{1}(M)$.
We say that $H$ is a virtual surface fibre subgroup of $\pi_{1}(M)$ if there exists some finite sheeted cover $M^{\prime}$ of $M$ whose fundamental group contains $H$ as a fibre subgroup. A finite sheeted covering of $M$ corresponds to a finite index subgroup of $\pi_{1}(M)$, so $\pi_{1}(M^{\prime})$ will be a finite index subgroup of $\pi_{1}(M)$ of which $H$ is a surface fibre subgroup.
@@
[undefal, Theorems 1, 2] Let $M$ be a compact and irreducible 3-manifold and let $H\unlhd\pi_{1}(M)$ be a finitely generated normal subgroup. If
$\pi_{1}(M)/H\cong\mathbb{Z}\text{ and }H\not\cong\mathbb{Z}/2\mathbb{Z},$
then $M$ can be expressed as a fibred 3-manifold with fibre surface $S$. Furthermore, we can choose $S$ such that $\pi_{1}(S)=H$
@@
[undefai, Theorem 2.3] Let $M$ be a closed and orientable irreducible 3-manifold, and suppose $\pi_{1}(M)\cong A*_{C}B$ where $C\neq A$ or $B$ and $C$ is isomorphic to the fundamental group of a closed surface $S$. Then there is an incompressible embedding of $S$ in $M$ separating $M$ into $M_{1}$ and $M_{2}$ with $\pi_{1}(M_{1})=A$, $\pi_{1}(M_{2})=B$ and $\pi_{1}(S)=C$.
More generally, the following powerful result essentially due to Stallings, Epstein and Waldhausen can be used to construct 2-sided incompressible surfaces from any graph of groups decomposition of the fundamental group of a closed and orientable 3-manifold. We will use the version that appears in [undefg].
@@
[undefg, Proposition 2.3.1] Let $M$ be a compact orientable 3-manifold. For any non-trivial graph of groups decomposition $(\Gamma,\mathfrak{G})$ of $\pi_{1}(M)$ with finitely many edge groups there exists a non-empty system $\overline{\Sigma}=\{\Sigma_{1},...,\Sigma_{n}\}$ of compact $2$-sided incompressible surfaces embedded in $M$, none of which are boundary parallel, such that for all $i$, $\operatorname{Im}(\pi_{1}(\Sigma_{i})\rightarrow\pi_{1}(M))$ is contained in some edge group of $(\Gamma,\mathfrak{G})$ and for all connected components $M_{j}$ of $M\backslash\overline{\Sigma}$, $\operatorname{Im}(\pi_{1}(M_{j})\rightarrow\pi_{1}(M))$ is contained in some vertex group of $(\Gamma,\mathfrak{G})$.
@@
In the paper of Culler and Shalen, the authors do not conclude that the surfaces they construct are compact. However, compactness is an immediate consequence of their proof — the surfaces are constructed as the connected components of the continuous preimage in the compact manifold of a closed set, so are themselves compact.
@@
[undefa, (K.18)] Let $M$ be a compact and orientable hyperbolic $3$-manifold with empty or toroidal boundary. Then $\pi_{1}(M)$ is hyperbolic relative to the conjugacy classes of the subgroups arising from the boundary components of $M$.
@@
Let $M$ be a compact and orientable hyperbolic $3$-manifold with empty or toroidal boundary. Then $M$ is irreducible.
@@
Let $M$ be as in the statement. Then the universal cover of the interior of $M$ is a copy of $\mathbb{H}^{3}$, which has vanishing second homotopy group as it is contractible. It follows that the second homotopy group of the interior of $M$ must also vanish, and the boundary of $M$ contains no spherical components, so $M$ is irreducible as required as it can contain no non-trivial sphere. ∎
@@
A compact and orientable hyperbolic $3$-manifold $M$ with empty or toroidal boundary that contains an embedded $2$-sided incompressible disc must be a solid torus. In particular, as a consequence of Theorem 4.8, $\pi_{1}(M)$ admits a non-trivial splitting over a trivial subgroup if and only if $M$ is a solid torus.
@@
For the first part of this corollary, assume that $M$ contains an embedded disc , so there exists an embedding of a disc $D$ into $M$ such that the boundary of $D$ is a non-trivial curve embedded in some boundary component $T$ of $M$. We will argue as in [undefq, Page 14, (3)] that $M$ is a solid torus.
@@
[undefa, (C.2)] Let $M$ be an orientable, irreducible $3$-manifold with empty or toroidal boundary and infinite fundamental group. Then $\pi_{1}(M)$ is torsion free.
Many of our methods will rely on a powerful result known as subgroup tameness to classify subgroups of the fundamental group of hyperbolic $3$-manifolds. For a full definition of the concept of geometric finiteness we direct the reader to the book 3-Manifold Groups by Aschenbrenner, Friedl and Wilton [undefa, Chapter 5]. For our purposes however, we will be able to define geometrically finite subgroups simply as relatively quasi-convex subgroups of hyperbolic $3$-manifold groups using the following theorem of Hruska.
@@
[undefs, Corollary 1.6] Let $G=\pi_{1}(M)$ be the fundamental group of a compact and orientable hyperbolic 3-manifold with empty or toroidal boundary and let $H$ be a finitely generated subgroup of G. Then $H$ is geometrically finite in $G$ if and only if $H$ is relatively quasi-convex with respect to the subgroups of $G$ that correspond to the fundamental groups of the boundary components.
@@
[undefa, Theorem 5.2] Let $M$ be a hyperbolic 3-manifold and let $H\leq\pi_{1}(M)$ be a finitely generated subgroup. Then either